Circular No. 8221

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

DE CIRCINI = POSSIBLE NOVA IN CIRCINUS W. Liller, Viña del Mar, Chile, writes that an objective-transmission-

diffraction-grating CCD spectrum (75 grooves/mm; aperture 100×100 mm; 0.20-m Schmidt camera), obtained on Oct. 11.002 UT, shows an exceedingly strong and broad Hα emission line (FWHM at least 5100 km/s); the brightness of its flat maximum was ≈ 7.7 times that of the surrounding continuum. A spectrogram obtained on Oct. 12.012 shows the Hα emission line basically unchanged, with its maximum brightness now estimated to be 6.2 times that of the surrounding continuum (FWHM = 5035 km/s). Photometry by Liller: Oct. 10.991, V = 10.27 ± 0.03, B − V = +0.39 ± 0.05; 12.001, V = 10.60 ± 0.03. L. A. G. Monard, Pretoria, S. Africa, reports that a CCD image taken on Oct. 11.704 yields mag 9.5 and the following precise position: α = 15^{h}17^{m}52^{s}.48, δ = −61^{o}57′16′′.4 (equinox 2000.0).

N. N. Samus, Institute of Astronomy, Moscow, informs us that the designation DE Cir has been given to this nova (cf. IAUC 8219).

SUPERNOVA 2003ir IN UGC 3726 Further to IAUC 8214, M. Ganeshalingam and W. Li report the LO-

TOSS discovery of an apparent supernova (mag ∼ 18.0) on unfiltered KAIT images taken on Oct. 11.5 and 12.5 UT. SN 2003ir is located at α = 7^{h}11^{m}08^{s}.18, δ = +25^{o}54′55′′.0 (equinox 2000.0), which is 15′′.2 east and 0′′.5 south of the nucleus of UGC 3726. A KAIT image taken on Apr. 8.2 showed nothing at this position (limiting mag ∼ 19.0).

V838 MONOCEROTIS D. K. Lynch and R. W. Russell, The Aerospace Corporation; and E.

Polomski, University of Minnesota, report 3–14 μm spectroscopy of V838 Mon on Oct. 11.6 and 12.6 UT at the Infrared Telescope Facility (+ BASS). The object showed a smooth, monotonically decreasing spectrum toward longer wavelengths with possible weak maxima at 8.7 and 10.6 μm. The double-peaked feature reported in February (IAUC 8078) has vanished, although there was evidence for a lingering absorption at 10.3 μm. The spectrum was much broader than can be represented by a Planck function, more closely resembling a power-law spectrum. Current infrared magnitudes: L = 3.52 ± 0.01, M = 2.70 ± 0.04, and N (10.5 μm) = 0.67 ± 0.02. Since February, the object’s brightness appears unchanged at M but has brightened by about half a magnitude at N.

2003 October 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green